HETAV PATEL

Grade 11,

Given: sec theta + tan theta =3, where theta belongs to (0,π/2). Find cos theta.

Given: sec theta + tan theta =3, where theta belongs to (0,π/2). Find cos theta.

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
As we know
sec²@ - tan²@ = 1
(Sec@ + tan@)(sec@ - tan@) = 1
Hence sec@ - tan@ = 1/3
Now
2sec @ = 3 + 1/3 = 10/3
sec @ = 5/3
Hence cos @ = 3/5